OpenStudy (anonymous):

a grasshopper jumps off the ground with an initial vertical velocity of 4 feet per second. After how many seconds does he reach the ground? please use the vertical motion formula h=-16t^2+vt+s

OpenStudy (anonymous):

someone help me on mine please

OpenStudy (cwrw238):

when he reaches ground h = 0 so 0 = -16t^2 + 4t + 0 t(16t - 4) = 0 t = 4/16 = 1/4 answer 1/4 second